OpenStudy (anonymous):

Using de Moivre’s theorem, express cos 6θ and sin 6θ in powers of cos θ and sin θ. Do i have to use binomial expansion??

OpenStudy (dumbcow):

yes i believe you do have to use binomial expansion \[\cos(6 \theta)+\sin(6 \theta) = (\cos \theta +\sin \theta)^{6}\]

OpenStudy (anonymous):

but it says using demoivre..so binomial isnt needed i think'

OpenStudy (anonymous):

isn't needed?? i don't understand..

OpenStudy (dumbcow):

wait i forgot the "i" the equality only holds if dealing with complex numbers \[\cos(6 \theta)+i \sin(6 \theta) = (\cos \theta +i \sin \theta)^{6}\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i always did like what @dumbcow did.. is there other alternative??

OpenStudy (anonymous):

if it says without demoivre,,then there are other wayss

OpenStudy (dumbcow):

\[=\cos^{6}+6i \cos^{5}\sin-15\cos^{4}\sin^{2}-20i \cos^{3}\sin^{3}+15\cos^{2}\sin^{4}+6i \cos \sin^{5}-\sin^{6}\]

OpenStudy (anonymous):

i think you dont need the expansion,as it said find in terms of powers of costheta

OpenStudy (anonymous):

find cos6theta in terms of costheta and sin6theta in terms of sintheta,are these separate or same question?

OpenStudy (anonymous):

find cos6θ and sin 6θ in powers of cos θ and sin θ.. the same.. i think i have to use the binomial..

OpenStudy (anonymous):

so i just have to expand it and that's it?? Do i have to separate it between the real and the imaginary in the end??

OpenStudy (anonymous):

the question is find cos6theta and sin6theta , it indicates they are different sums........ it is "and", why take it as +

OpenStudy (dumbcow):

soham, because de'moivres theorem is mentioned i think it was implied mandy, yes i believe you would have to separate real and imaginary coefficients

OpenStudy (anonymous):

Thanks a lot!! Im not lost anymore.. ^^

OpenStudy (anonymous):

right